Obtain a standard deck of 52 cards, and think of the cards as the 52 items your firm produces each day. Let the four aces and four kings in the deck represent defective items.
if 10 items are randomly sampled from a days production, what should be the proportion of defective items?
Please help! Thank you!

Answer by stanbon(75887)

Obtain a standard deck of 52 cards, and think of the cards as the 52 items your firm produces each day. Let the four aces and four kings in the deck represent defective items.
That is 8 defective in 52
------
If 10 items are randomly sampled from a days production, what should be the proportion of defective items?
------
Use a proportion:
x/10 = 8/52
---
x = 10(8/52)
---
x = 10(2/13)
---
x = 20/13 = 1.54 items
===============================
Cheers,
Stan H.